SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Mar.28, 2006)

Attendees: Rainer Bartoldus, Al Eisner, Fang Fang, Marc Kelly, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker

Operations

  • We have been taking cosmics runs because of the downtime. A new version of the TSF firmware has been installed since March 27 swing shift (See the DCT/DCZ section below for more details).

EMT

  • The online EMT ROM lib was updated and tested with cosmics after the TRG operation meeting today.

DCT/DCZ

  • Marc got a new and improved version of the TSF firmware working at IR2 and the test stand yesterday. The firmware issue finally got resolved. See HN posted. Marc gave us a brief summary and status report on the firmware changes made. Marc will check in the firmware code to the CVS for the record in case we ever need it again in the future.

GLT

Online

  • Fang is working on shipping the input datagram which was tested at the test stand but got segmentation errors. Rainer will discuss with Fang concerning this matter after the meeting.

L3

  • Arik and Rainer are working on implementation of the improved L3 tracking code. They are currently running on the L1passthru run from the end of last year to predict its impact on hadron ratio, lumi, etc. The code will get impletemented after we get smooth data-taking condition.

Data Quality and Monitoring

  • Rahmat is working on updating some of the instructions/information for DQM in the Trigger JAS pages.
  • We talk about creating strip charts for OPR monitoring by using ratio of certain (L3 output) trigger lines that relate to luminosity monitoring.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.